OpenStudy (anonymous):

Which literal equation is not equivalent to x = 2y + z ? x - 2y = z x-z=2y y = x-z/2 x + 2y = z

OpenStudy (anonymous):

x - 2y = z -- add 2y to each side x = 2y + z this is equivalent...so its not this one x - z = 2y -- add z to both sides x = 2y + z this is equivalent....not this one y = (x - z) / 2 -- multiply by 2 2y = x - z -- add z to both sides 2y + z = x or x = 2y + z this is equivalent...not this one x + 2y = z -- subtract 2y from both sides x = -2y + z this is NOT equivalent.....IT IS THIS ONE